OpenStudy (anonymous):

Is 4x-3=19 true false or open

OpenStudy (anonymous):

@Ashleyisakitty I was Gonna Say Open Since We dont know what x is

OpenStudy (anonymous):

Is that Right

OpenStudy (anonymous):

4x-3=19 First solve for x Then substitute aferwards, do you know how to do that

OpenStudy (anonymous):

Yes Thanks

OpenStudy (anonymous):

lol np, if u have any further questions, let me no :)

OpenStudy (anonymous):

but how do I know what x is

OpenStudy (anonymous):

solve for x i'll show it to u if ud like

OpenStudy (anonymous):

Yes Thnaks I Would Like That

OpenStudy (anonymous):

Typo lol Thanks

OpenStudy (anonymous):

Its True 4x-3=9 4x=19+3 4x=22 x=22/4 x=5.5 means 4*5.5-3=19 is true

OpenStudy (anonymous):

4x-3=19 you must isolate the x, so add three on both sides cuz its a negitive if it was a positive it would be the other way around 4x=19+3 4x=22 Now divide by 4 on both sides x=5.5

OpenStudy (anonymous):

is that correct?

OpenStudy (anonymous):

Yes Thanks Guys

OpenStudy (anonymous):

Thanks alot

OpenStudy (anonymous):

yupeee....you are welcome. :)

OpenStudy (anonymous):

@callie2240 can i tag you if i have more questions

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Kk thanks allot @callie2240 and @vaibhavm

OpenStudy (anonymous):

np

OpenStudy (anonymous):

gtg for now though, dinner

OpenStudy (anonymous):

Oh Okay

OpenStudy (anonymous):

back

OpenStudy (anonymous):

haha okay good

OpenStudy (anonymous):

lol ya so anymore questions

OpenStudy (anonymous):

we are ready.......lol

OpenStudy (anonymous):

Haha Okay Yeah Ill Post Them :)

OpenStudy (anonymous):

aight, let me know which one lol

OpenStudy (anonymous):

it seems callie desperate to answer.......lol